### v1

@@
The inventory routing problem (IRP) focuses on jointly optimizing inventory and distribution operations from a supplier to retailers over multiple days. Compared to other problems from the vehicle routing family, the interrelations between inventory and routing decisions render IRP optimization more challenging and call for advanced solution techniques. A few studies have focused on developing large neighborhood search approaches for this class of problems, but this remains a research area with vast possibilities due to the challenges related to the integration of inventory and routing decisions. In this study, we advance this research area by developing a new large neighborhood search operator tailored for the IRP. Specifically, the operator optimally removes and reinserts all visits to a specific retailer while minimizing routing and inventory costs. We propose an efficient tailored dynamic programming algorithm that exploits preprocessing and acceleration strategies. The operator is used to build an effective local search routine, and included in a state-of-the-art routing algorithm, i.e., Hybrid Genetic Search (HGS). Through extensive computational experiments, we demonstrate that the resulting heuristic algorithm leads to solutions of unmatched quality up to this date, especially on large-scale benchmark instances.
@@ 1 Introduction
The use of large neighborhood search (LNS) techniques has permitted significant progress in the solution of various vehicle routing problems (VRPs) (see, e.g., Christiaens and Vanden Berghe 2020, Pacheco et al. 2023b) and combinatorial optimization (CO) problems in general (Hendel 2022). Especially when multiple decisions are interrelated, classical local search neighborhoods often do not allow an efficient search space exploration, while larger neighborhoods can be instrumental in exploring structurally different solutions.
The Inventory Routing Problem (IRP), in particular, is a prototypical example of a problem involving multiple interrelated decisions. In its canonical form, the IRP requires jointly optimizing inventory and delivery routes from a supplier to geographically dispersed retailers over a planning horizon measured in days. This involves a complex set of decisions that include the daily selection of retailers receiving visits, their delivery quantities, the assignment of vehicles to these visits, and their ordering. Despite a few recent and promising works on the topic (see, e.g., Chitsaz et al. 2019, Diniz et al. 2020, Archetti et al. 2021, Skålnes et al. 2023), applications of large neighborhood search for this problem remain few and far. Many large neighborhood searches designed for IRP fundamentally rely on mixed integer programming (see, e.g., Bouvier et al. 2024a), which can become time-consuming and not permit many iterations. For this reason, though extensive progress has been made on solution methods over the past decades, judging from the average gap and standard deviation of state-of-the-art methods, significant improvements are still possible on medium and larger-scale instances such as the ones introduced in Skålnes et al. (2024).
@@
    We propose a new large neighborhood operator that removes and reinserts all visits to a specific retailer while minimizing routing and inventory costs. While this may evoke the concepts used in ruin-and-recreate methods (Ropke and Pisinger 2006), the way we apply this strategy is fundamentally different. We demonstrate that, after preprocessing, the reinsertion can be done optimally using efficient dynamic programming and acceleration strategies specifically tailored to this case. Moreover, the speed of our method permits us to explore this large neighborhood fully for all retailers in random order until a local minimum is reached.
@@
    We integrate this LNS operator into the hybrid genetic search (HGS) of (Vidal et al. 2012, Vidal 2022), extending this method to solve IRPs. Specifically, the new operator replaces the neighborhood search operator originally used in (Vidal et al. 2012) to choose the allowed visit-day combinations by inspection in the context of multi-period VRP. This is the first time an operator has been designed for an IRP that optimizes replenishment decisions across the entire planning horizon.
@@
    We extend the IRP to the Stock-Out Allowed (SOA) model by introducing a penalty term for stock-out. This variant has been rarely considered, but it is especially relevant when applying IRP solution techniques in a dynamic environment, e.g., as studied in (Greif et al. 2024), as this often requires the ability to evaluate infeasible solutions in some scenarios.
@@
    We conduct an extensive computational campaign to evaluate the neighborhood search efficiency and compare the adapted HGS’s performance with other state-of-the-art IRP algorithms. Most notably, the resulting method identifies 187 new best solutions out of 800 small data instances and 178 new best solutions in 240 large instances. Even more remarkable is its performance when applied to the new challenging IRP dataset proposed in Skålnes et al. (2024), where it finds 191 new benchmark solutions on 270 instances with a shorter average running time. These findings highlight the efficiency of the proposed enhanced HGS algorithm, thanks to the development of the new large neighborhood operator.
@@ 2 Related works
Ahuja et al. (2002) reviews various large neighborhood search techniques, categorizing them into three main categories. The first category, variable-depth methods, employs heuristics to explore increasingly large (higher-order) neighborhoods. The heuristic of Lin and Kernighan (1973) is a seminal example, as it effectively harnesses pruning techniques to explore many solutions obtained by replacing a variable number of edges. Applications of this technique to various CO problems are demonstrated in Applegate et al. (2003) and Hendel (2022). The second category includes network flow-based improvement algorithms, which rely on solving network flow subproblems to explore large neighborhoods. Ejection chains (Thompson and Psaraftis 1993, Vidal 2017) and assignment-based neighborhoods (Toth and Tramontani 2008, Capua et al. 2018) are notable examples, leveraging shortest path and assignment problems to explore an exponential subset of solutions in polynomial time. The third category, neighborhoods induced by restrictions, exploits special cases of the original problem that admit polynomial-time solutions. For example, certain restrictions of the Traveling Salesman Problem (TSP) can be solved in polynomial time, leading to various classes of large neighborhoods that are broadly applicable to routing problems (Balas and Simonetti 2001, Pacheco et al. 2023a). Additionally, complex CO problems can often be decomposed into several classes of decision variables, allowing for local searches on a subset of these variables. When combined with a move evaluation procedure that optimally determines the remaining variables, this approach can lead to effective searches in large neighborhoods (Vidal 2017, Toffolo et al. 2019). Collectively, these techniques have led to important breakthroughs for difficult problems.
@@
The Inventory Routing Problem (IRP) is a foundational problem in supply chain optimization that has been extensively studied over the past decades, as highlighted in surveys by Andersson et al. (2010), Coelho et al. (2014), and Vidal et al. (2020). The IRP is a canonical problem designed to capture the challenges associated with the simultaneous optimization of both inventory replenishment strategies and routing plans, offering significant potential for cost reduction in supply chain operations.
@@
The origins of the IRP trace back to the work of Bell et al. (1983), who introduced the concept for a gas delivery application. Their approach leveraged Lagrangian relaxation to decompose the problem across temporal and vehicle dimensions. Later, Christiansen (1999) expanded the IRP’s application to maritime logistics, coining the term “inventory pickup and delivery problem.” Their approach used Dantzig-Wolfe decomposition and column-generation techniques. Further advancements in addressing the IRP were made by Carter et al. (1996) and Campbell and Savelsbergh (2004), who developed efficient heuristic methods by decomposing the IRP into an allocation problem and a routing problem.
@@
Numerous metaheuristic strategies have since been proposed for the IRP. Examples include tabu search (Rusdiansyah and Tsao 2005), genetic algorithms (Abdelmaguid and Dessouky 2006), a greedy randomized adaptive search procedure (Campbell and Savelsbergh 2004), adaptive large neighborhood search (LNS) (Coelho et al. 2012), and a variable neighborhood descent heuristic with stochastic neighborhood ordering (Alvarez et al. 2018), which combines routing-focused and joint routing-inventory operators. In addition, “matheuristics”, which integrate heuristic search with mathematical programming, have shown significant promise. Notable examples include hybrid methods by Archetti et al. (2012, 2017), Chitsaz et al. (2019), Archetti et al. (2021), Skålnes et al. (2023), Bouvier et al. (2024b) and Dinh et al. (2025). Among these, Skålnes et al. (2023) developed a matheuristic that is invoked on each primal solution identified by a branch-and-cut algorithm. Whenever the mathematical program discovers a new best upper bound, this method attempts to further improve it by applying mixed integer programming to a restricted problem, which involves removing and reinserting clusters of clients. This advanced approach successfully identified 741 best-known solutions out of 878 existing benchmark instances and proved the optimality of 458 of these solutions.
@@
Research has also extended to various IRP variants. For instance, Gaur and Fisher (2004) addressed the periodic IRP, optimizing replenishment under cyclical demand, while other studies focused on cyclic (Gunawan et al. 2019, Vincent et al. 2021), multi-product (Ohmori and Yoshimoto 2021, Mahjoob et al. 2022), and multi-depot (Guimarães et al. 2019) IRPs. Strategic extensions integrating production decisions have also been explored (Adulyasak et al. 2014b). As noted by Andersson et al. (2010), early research predominantly relied on heuristics due to the problem’s complexity. However, subsequent works introduced exact algorithms, primarily branch-and-cut methods (Coelho and Laporte 2013a, b, Adulyasak et al. 2014a, Archetti et al. 2014, Coelho and Laporte 2014), with one notable exception employing branch-and-price (Desaulniers et al. 2016).
@@
Despite their efficacy for small to medium-scale problems, these methods often struggle with scalability, particularly for instances involving more than 50 retailers or planning horizons exceeding three days. Addressing this gap, Skålnes et al. (2024) proposed challenging large-scale IRP instances to stimulate further research. Recently, Bouvier et al. (2024b) introduced a large neighborhood search procedure designed for very large IRPs involving an average of 30 commodities, 16 depots, and 600 retailers. However, the scale of the problems they address and the complexity of the underlying neighborhoods limit the procedure to only a handful of local search descents. This highlights the ongoing need for efficient neighborhood operators —both local and large— that can effectively tackle large-scale IRP instances.
@@ 3 Problem Statement
The IRP is defined on a complete directed graph $\mathcal{G}=(\mathcal{N},\mathcal{A})$, where the node set $\mathcal{N}=\{0,n+1\}\cup\mathcal{N}^{\prime}$ includes the departure depot at node 00, the destination depot at node $n+1$, and the retailer nodes $\mathcal{N}^{\prime}=\{1,\ldots,n\}$. The arc set $\mathcal{A}$ represents paths between nodes, with each arc $(i,j)\in\mathcal{A}$ associated with a non-negative cost $c_{i,j}$. The planning horizon spans $H$ days, denoted by $\mathcal{T}=\{1,\ldots,H\}$. Each retailer node $i\in\mathcal{N}^{\prime}$ has a per-day inventory holding cost $h_{i}$, an initial inventory level $I_{i}^{0}$, and a maximum inventory level $U_{i}$. In the Stock-Out Allowed (SOA) model, $\rho h_{i}$ represents the per-unit stock-out penalty for retailer $i$, with $\rho>1$. In contrast, in the No Stock-Out (NSO) model, the value of $\rho$ is set to infinity. Each day $t\in\mathcal{T}$, $d_{0}^{t}$ units of product are made available at the supplier, and $d_{i}^{t}$ units are consumed by retailer $i\in\mathcal{N}^{\prime}$. A fleet $\mathcal{K}$ of $K$ identical vehicles, each with capacity $Q$, provides the delivery service.
The IRP involves creating a distribution plan that minimizes the total cost. This plan specifies, for each day $t\in\mathcal{T}$, the delivery quantity to each retailer $i\in\mathcal{N}^{\prime}$ and the routes of vehicles for servicing the retailers on that day. The total cost encompasses the holding costs at both the supplier and retailers, the stock-out costs under the SOA model, and the transportation costs associated with the routes.
@@
A mathematical formulation of the IRP under the SOA model is presented below, extending the formulation from Archetti et al. (2007) to account for possible stock-outs. Let $I_{i}^{t}$ represent the inventory level at retailer $i$ at the end of the day $t$, and $B_{i}^{t}$ the unsatisfied demand on that day. Also, $I_{0}^{t}$ is the inventory level at the supplier at the end of the day $t$. Binary variable $z^{k,t}_{i}$ is set to one if node $i\in\mathcal{N}$ is visited by vehicle $k\in\mathcal{K}$ on day $t\in\mathcal{T}$. Continuous variable $q^{k,t}_{i}$ denotes the quantity delivered to retailer $i\in\mathcal{N^{\prime}}$ on day $t$ by vehicle $k$. Binary variable $y^{k,t}_{i,j}$ indicates whether vehicle $k$ traverses arc $(i,j)\in\mathcal{A}$ on day $t$. Finally, $Q_{i}^{k,t}$ is a continuous variable representing the load of vehicle $k$ upon arrival at node $i$, reflecting the cumulative load at that point in the vehicle’s route. The objective is to minimize the total cost, which includes holding costs at the supplier $h_{0}I_{0}^{t}$, holding costs at the retailers $\sum_{i\in\mathcal{N^{\prime}}}h_{i}I_{i}^{t}$, penalty costs for stock-outs $\sum_{i\in\mathcal{N^{\prime}}}\rho h_{i}B_{i}^{t}$, and transportation costs $\sum_{k\in\mathcal{K}}\sum_{(i,j)\in\mathcal{A}}c_{i,j}y^{k,t}_{i,j}$. The resulting mathematical formulation is:
@@
$Q_{0}^{k,t}=\sum_{\forall\,i\in\mathcal{N^{\prime}}}q_{i}^{k,t}\quad$ $k\in\mathcal{K},t\in\mathcal{T},$ (10)
@@
Property 1 (Archetti and Ljubić (2022))

### v2

@@
The inventory routing problem (IRP) focuses on jointly optimizing inventory and distribution operations from a supplier to retailers over multiple days. Compared to other problems from the vehicle routing family, the interrelations between inventory and routing decisions render IRP optimization more challenging and call for advanced solution techniques. A few studies have focused on developing large neighborhood search approaches for this class of problems, but this remains a research area with vast possibilities due to the challenges related to the integration of inventory and routing decisions. In this study, we advance this research area by developing a new large neighborhood search operator tailored for the IRP. Specifically, the operator optimally removes and reinserts all visits to a specific retailer while minimizing routing and inventory costs. We propose an efficient tailored dynamic programming algorithm that exploits preprocessing and acceleration strategies. The operator is used to build an effective local search routine, and included in a state-of-the-art routing algorithm, i.e., Hybrid Genetic Search. Through extensive computational experiments, we demonstrate that the resulting heuristic algorithm leads to solutions of unmatched quality up to this date, especially on large-scale benchmark instances.
@@ 1 Introduction
The use of large neighborhood search (LNS) techniques has permitted significant progress in the solution of various vehicle routing problems (VRPs) (see, e.g., christiaens2020slack, pacheco2023exponential) and combinatorial optimization (CO) problems in general (hendel2022adaptive). Especially when multiple decisions are interrelated, classical local search neighborhoods often do not allow an efficient search space exploration, while larger neighborhoods can be instrumental in exploring structurally different solutions.
The Inventory Routing Problem (IRP), in particular, is a prototypical example of a problem involving multiple interrelated decisions. In its canonical form, the IRP requires jointly optimizing inventory and delivery routes from a supplier to geographically dispersed retailers over a planning horizon measured in days. This involves a complex set of decisions that include the daily selection of retailers receiving visits, their delivery quantities, the assignment of vehicles to these visits, and their ordering. Despite a few recent and promising works on the topic (see, e.g., chitsaz2019unified, diniz2020efficient, archetti2021kernel, skaalnes2023branch), applications of LNS for this problem remain few and far. Many LNS designed for IRP fundamentally rely on mixed integer programming (see, e.g., bouvier2024solving), which can become time-consuming and not permit many iterations. For this reason, though extensive progress has been made on solution methods over the past decades, judging from the average gap and standard deviation of state-of-the-art methods, significant improvements are still possible on medium and larger-scale instances such as the ones introduced in skaalnes2024new.
@@
    We propose a new large neighborhood operator that removes and reinserts all visits to a specific retailer while minimizing routing and inventory costs. While this may evoke the concepts used in ruin-and-recreate methods (ropke2006adaptive), the way we apply this strategy is fundamentally different. We demonstrate that, after preprocessing, the reinsertion can be done optimally using efficient dynamic programming and acceleration strategies specifically tailored to this case. Moreover, the speed of our method permits us to explore this large neighborhood fully for all retailers in random order until a local minimum is reached.
@@
    We integrate this LNS operator into the hybrid genetic search (HGS) of (vidal2012hybrid, vidal2022hybrid), extending this method to solve IRPs. Specifically, the new operator replaces the neighborhood search operator originally used in vidal2012hybrid to choose the allowed visit-day combinations by inspection in the context of multi-period VRP. This is the first time an operator has been designed for an IRP that optimizes replenishment decisions across the entire planning horizon.
@@
    We extend the IRP to the Stock-Out Allowed (SOA) model by introducing a penalty term for stock-out. This variant has been rarely considered, but it is especially relevant when applying IRP solution techniques in a dynamic environment, e.g., as studied in (greif2024combinatorial), as this often requires the ability to evaluate infeasible solutions in some scenarios.
@@
    We conduct an extensive computational campaign to evaluate the neighborhood search efficiency and compare the adapted HGS’s performance with other state-of-the-art IRP algorithms. Most notably, the resulting method identifies 187 new best solutions out of 800 small data instances and 178 new best solutions in 240 large instances. Even more remarkable is its performance when applied to the new challenging IRP dataset proposed in skaalnes2024new, where it finds 191 new benchmark solutions on 270 instances with a shorter average running time. These findings highlight the efficiency of the proposed enhanced HGS algorithm, thanks to the development of the new large neighborhood operator.
@@ 2 Related works
ahuja2002survey reviews various large neighborhood search techniques, categorizing them into three main categories. The first category, variable-depth methods, employs heuristics to explore increasingly large (higher-order) neighborhoods. The heuristic of lin1973effective is a seminal example, as it effectively harnesses pruning techniques to explore many solutions obtained by replacing a variable number of edges. Applications of this technique to various CO problems are demonstrated in applegate2003traveling and hendel2022adaptive. The second category includes network flow-based improvement algorithms, which rely on solving network flow subproblems to explore large neighborhoods. Ejection chains (Thompson1993, Vidal2017b) and assignment-based neighborhoods (Toth2008, Capua2018) are notable examples, leveraging shortest path and assignment problems to explore an exponential subset of solutions in polynomial time. The third category, neighborhoods induced by restrictions, exploits special cases of the original problem that admit polynomial-time solutions. For example, certain restrictions of the Traveling Salesman Problem (TSP) can be solved in polynomial time, leading to various classes of large neighborhoods that are broadly applicable to routing problems (Balas2001, pacheco2023exponential). Additionally, complex CO problems can often be decomposed into several classes of decision variables, allowing for local searches on a subset of these variables. When combined with a move evaluation procedure that optimally determines the remaining variables, this approach can lead to effective searches in large neighborhoods (Vidal2017b, Toffolo2019). Collectively, these techniques have led to important breakthroughs for difficult problems.
@@
The Inventory Routing Problem (IRP) is a foundational problem in supply chain optimization that has been extensively studied over the past decades, as highlighted in surveys by anderssoN2010industrial, coelho2014survey and vidal2020survey. The IRP is a canonical problem designed to capture the challenges associated with the simultaneous optimization of both inventory replenishment strategies and routing plans, offering significant potential for cost reduction in supply chain operations.
@@
The origins of the IRP trace back to the work of bell1983improving, who introduced the concept for a gas delivery application. Their approach leveraged Lagrangian relaxation to decompose the problem across temporal and vehicle dimensions. Later, christianseN1999decomposition expanded the IRP’s application to maritime logistics, coining the term “inventory pickup and delivery problem.” Their approach used Dantzig-Wolfe decomposition and column-generation techniques. Further advancements in addressing the IRP were made by carter1996solving and campbell2004decomposition, who developed efficient heuristic methods by decomposing the IRP into an allocation problem and a routing problem.
@@
Numerous metaheuristic strategies have since been proposed for the IRP. Examples include tabu search (rusdiansyah2005integrated), genetic algorithms (abdelmaguid2006genetic), a greedy randomized adaptive search procedure (campbell2004decomposition), adaptive large neighborhood search (LNS) (coelho2012inventory), and a variable neighborhood descent heuristic with stochastic neighborhood ordering (alvarez2018iterated), which combines routing-focused and joint routing-inventory operators. In addition, “matheuristics”, which integrate heuristic search with mathematical programming, have shown significant promise. Notable examples include hybrid methods by archetti2012hybrid, archetti2017matheuristic, chitsaz2019unified, archetti2021kernel, skaalnes2023branch, Bouvier2024 and dinh2025matheuristic. Among these, skaalnes2023branch developed a matheuristic that is invoked on each primal solution identified by a branch-and-cut algorithm. Whenever the mathematical program discovers a new best upper bound, this method attempts to further improve it by applying mixed integer programming to a restricted problem, which involves removing and reinserting clusters of clients. This advanced approach successfully identified 741 best-known solutions out of 878 existing benchmark instances and proved the optimality of 458 of these solutions.
@@
Research has also extended to various IRP variants. For instance, gaur2004periodic addressed the periodic IRP, optimizing replenishment under cyclical demand, while other studies focused on cyclic (gunawaN2019simulated, vincent2021multi), multi-product (ohmori2021multi, mahjoob2022modified), multi-depot (guimaraes2019two), multi-attribute (coelho2020variable), and bi-objective (feng2025bi) IRPs. Strategic extensions integrating production decisions have also been explored (adulyasak2014optimization) As noted by anderssoN2010industrial, early research predominantly relied on heuristics due to the problem’s complexity. However, subsequent works introduced exact algorithms, primarily branch-and-cut methods (CoelhoLaporte2013, CoelhoLaporte2013b, adulyasak2014formulations, Archetti2014, CoelhoLaporte2014), with one notable exception employing branch-and-price (DesaulniersRC2016). Other contributions include VND methods (larrain2017variable) and matheuristic frameworks (vincent2025three) that combine mathematical programming with metaheuristics to address larger-scale instances.
@@
Despite their efficiency for small to medium-scale problems, these methods often struggle with scalability, particularly for instances involving more than 50 retailers or planning horizons exceeding three days. Addressing this gap, skaalnes2024new proposed challenging large-scale IRP instances to stimulate further research. Recently, Bouvier2024 introduced a large neighborhood search procedure designed for very large IRPs involving an average of 30 commodities, 16 depots, and 600 retailers. However, the scale of the problems they address and the complexity of the underlying neighborhoods limit the procedure to only a handful of local search descents. This highlights the ongoing need for efficient neighborhood operators —both local and large— that can effectively tackle large-scale IRP instances.
@@ 3 Problem Statement
Our IRP is defined on a complete directed graph $\mathcal{G}=(\mathcal{N},\mathcal{A})$, where the node set $\mathcal{N}=\{0,n+1\}\cup\mathcal{N}^{\prime}$ includes the departure depot at node 0, the destination depot at node $n+1$, and the retailer nodes $\mathcal{N}^{\prime}=\{1,\ldots,n\}$. The arc set $\mathcal{A}$ represents paths between nodes, with each arc $(i,j)\in\mathcal{A}$ associated with a non-negative cost $c_{i,j}$. The planning horizon spans $H$ days, denoted by $\mathcal{T}=\{1,\ldots,H\}$. Each retailer node $i\in\mathcal{N}^{\prime}$ has a per-day inventory holding cost $h_{i}$, an initial inventory level $I_{i}^{0}$, and a maximum inventory level $U_{i}$. In the Stock-Out Allowed (SOA) model, $\rho h_{i}$ represents the per-unit stock-out penalty for retailer $i$, with $\rho>1$. In contrast, in the No Stock-Out (NSO) model, the value of $\rho$ is set to infinity. Each day $t\in\mathcal{T}$, $d_{0}^{t}$ units of product are made available at the supplier, and $d_{i}^{t}$ units are consumed by retailer $i\in\mathcal{N}^{\prime}$. A fleet $\mathcal{K}$ of $K$ identical vehicles, each with capacity $Q$, provides the delivery service.
The IRP involves creating a distribution plan that minimizes the total cost. This plan specifies, for each day $t\in\mathcal{T}$, the delivery quantity to each retailer $i\in\mathcal{N}^{\prime}$ and the routes of vehicles for servicing the retailers on that day. The total cost encompasses the holding costs at both the supplier and retailers, the stock-out costs under the SOA model, and the transportation costs associated with the routes. In the IRP, allowing stock-outs is a realistic modeling approach that reflects many real-world scenarios. This no-backlogging setup is particularly suited to situations where unmet demand cannot or should not be fulfilled later. For example, in fast-moving consumer goods, customers typically turn to competitors if products are out of stock, meaning lost sales rather than deferred ones. Similarly, in retail, pharmaceuticals, or seasonal product delivery, stock-outs represent immediate service failure, and backlogging is irrelevant or impractical. The penalty cost serves as a proxy for lost revenue, customer dissatisfaction, or contractual penalties, and it incentivizes timely and adequate deliveries. This modeling choice also simplifies the mathematical formulation by avoiding the complexity of backlog tracking while maintaining economic pressure to avoid stock-outs.
@@
A mathematical formulation of the IRP under the SOA model is presented below, extending the formulation from archetti2007branch to account for possible stock-outs. Let $I_{i}^{t}$ represent the inventory level at retailer $i$ at the end of the day $t$, and $B_{i}^{t}$ the unsatisfied demand on that day. Also, $I_{0}^{t}$ is the inventory level at the supplier at the end of the day $t$. Binary variable $z^{k,t}_{i}$ is set to one if node $i\in\mathcal{N}$ is visited by vehicle $k\in\mathcal{K}$ on day $t\in\mathcal{T}$. Continuous variable $q^{k,t}_{i}$ denotes the quantity delivered to retailer $i\in\mathcal{N^{\prime}}$ on day $t$ by vehicle $k$. Binary variable $y^{k,t}_{i,j}$ indicates whether vehicle $k$ traverses arc $(i,j)\in\mathcal{A}$ on day $t$. Finally, $Q_{i}^{k,t}$ is a continuous variable representing the load of vehicle $k$ upon arrival at node $i$, reflecting the cumulative load at that point in the vehicle’s route. The objective is to minimize the total cost, which includes holding costs at the supplier $h_{0}I_{0}^{t}$, holding costs at the retailers $\sum_{i\in\mathcal{N^{\prime}}}h_{i}I_{i}^{t}$, penalty costs for stock-outs $\sum_{i\in\mathcal{N^{\prime}}}\rho h_{i}B_{i}^{t}$, and transportation costs $\sum_{k\in\mathcal{K}}\sum_{(i,j)\in\mathcal{A}}c_{i,j}y^{k,t}_{i,j}$. The resulting mathematical formulation is:
@@
$Q_{0}^{k,t}=\sum_{i\in\mathcal{N^{\prime}}}q_{i}^{k,t}\quad$ $k\in\mathcal{K},t\in\mathcal{T},$ (10)
@@
Property 1 (archetti2022comparison)